OpenStudy (anonymous):

FACTOR X^2 + 6x + 8 = 0

OpenStudy (anonymous):

4 + 2 = 6 4 x 2 = 8 so maybe the factors are ....?

OpenStudy (anonymous):

those arent the factors @mebs ?

OpenStudy (anonymous):

They are but how do you write it?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

lets see I'll say the answer and you try this one ANS: (x+4)(x+2) try: x^2 + 4x + 4

OpenStudy (anonymous):

Find two numbers that add to 4 and multiply to 4.

OpenStudy (anonymous):

x^2 +6x +8=0 =x +4 x +2 =(x+4)(x+2)=0

OpenStudy (anonymous):

is it...isn't it?

OpenStudy (anonymous):

When you're factorising this you're basically looking for two brackets thats when you multiply them out get the expression above. Because you've got an x^2 you know at some stage you have to multiply x by x. So you know you're brackets will look something like (x+....)(x+....) Now to find the ..... values. You need two numbers that add up +6 and multiply to get +8 (if you multiply the brackets back out once we have a solution the reasons for this should become clear. So the only two numbers that satisfy those conditions are +4 and +2 (i.e. +2+6 = +8 and (+2)(+4)= +8) so x^2 + 6x + 8 = (x+4)(x+2) hope this helps!